OpenStudy (anonymous):

Find the derivative of the function f(x) = (x^2 - 1)(x^2 - 2).

hartnn (hartnn):

do you know about product rule ?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

that doesn't mean my class didn't learn it. I've been out for a while

OpenStudy (anonymous):

\((fg)' = f'g + fg'\)

hartnn (hartnn):

let u and v be the 2 functions of 'x' then, \(\large (uv)' = u'v+uv'\)

OpenStudy (luigi0210):

You could try distributing?

OpenStudy (anonymous):

It is less work to expand and just apply the power rule.

hartnn (hartnn):

here, take u= x^2-1 v=x^2-1 OR, as lui mentioned, you can distribute! in which case you won't need product rule at all!

OpenStudy (anonymous):

i am confused

OpenStudy (anonymous):

 This is what @Luigi0210 meant by distribution, now just find the derivative of this by using the power rule.